### v1

# Matrix Multiplication in the MPC Model

Abstract
In this paper, we study the matrix multiplication problem in the MPC model. We have two matrices, and the task is to compute their product. These matrices are evenly distributed over $P$ processors. Each processor has $M$ memory such that $P\cdot M\geq$ (size of the matrices). The computation proceeds in synchronous rounds. In a communication round, a processor can send and receive messages to(from) any other processor, with the total size of messages sent or received being $O(M)$. We give an almost complete characterisation of the problem in various settings. We prove tight upper bounds and lower bounds for the problems in three different settings–when the given input matrices are (i) general square matrices, (ii) rectangular matrices, and (iii) sparse square matrices (that is, each row and column contains a bounded number of nonzero elements). In particular, we prove the following results:
  1. 1.
    Multiplication of two $n\times n$ matrices in the MPC model with $n^{\alpha}$ processors each with $O(n^{2-\alpha})$ memory, requires $\Theta(n^{\frac{\alpha}{2}})$ rounds in semirings.
  2. 2.
    Multiplication of two rectangular matrices of size $n\times d$ and $d\times n$ (where $d\leq n$) respectively, with $n$ processors of $O(n)$ memory requires $\Theta(\frac{d}{\sqrt{n}})$ rounds in semirings.
  3. 3.
    Multiplication of two rectangular matrices of size $d\times n$ and $n\times d$ ( where $d\leq n$) respectively requires
    1. (a)
      $\Theta(\sqrt{d}+\log_{d}n)$ rounds with $n$ processors and $O(d)$ memory per processor in semirings
    2. (b)
      $\Theta(\frac{d}{\sqrt{n}})$ rounds with $d$ processors and $O(n)$ memory per processor in semirings.
  4. 4.
    Multiplication of two $d$-sparse matrices (each row and column of the matrices contains at most $d$-nonzero elements) with $n$ processors and $O(d)$ memory per processor can be done in $O(d^{0.9})$ rounds in semirings.
keywords:
Massively parallel computing, distributed algorithms, , matrix multiplication
category:
[…]
1 Introduction
We are given two matrices $A$ and $B$, and our task is to compute $C=A\times B$. The complexity to solve this problem differs with the dimensions and structure of the input matrices. Therefore, we explore the problem for the following cases that are often studied: when matrices $A$ and $B$ are (i) square matrices i.e. both of dimension $n\times n$, (ii) rectangular matrices of dimension $n\times d$ and $d\times n$, (iii) rectangular matrices of dimension $d\times n$ and $n\times d$, (iv) $d$- sparse matrices i.e. both are $n\times n$ size matrices such that each matrix contains most $d$ non-zero elements in each row and each column. The problem has been studied for all these cases in various models of distributed and parallel algorithms, for example, the clique model, the low bandwidth model etc. [6, 3, 2, 1, 4, 5]. We study this problem in the MPC model.
1.1 MPC Model
In models like clique and the low bandwidth, the primary restriction is put on the amount of communication between the nodes, whereas the internal computation within a node (or processor) remains unbounded in terms of time and space. In the MPC model, local memory is the primary bottleneck. Consider an input problem of size $N$. To solve the problem in the MPC model, we have $P$ processors, each with $M$ memory. The input is distributed among $P$ processors. Therefore, we have the following natural constraints: $P\cdot M\geq N$ and $P\cdot M\geq Output-size$ in order to store the complete input and output, respectively. The computation proceeds in synchronous rounds. Processors perform computation locally, and at the end of a round, they exchange their messages. Because of the memory restriction, each processor is allowed to send and receive $O(M)$ amount of information in each round. If any processor receives more than $O(M)$ amount of information during the course of any algorithm in the MPC model, the algorithm is considered failed.
1.2 Setup
In this paper, we primarily focus on the matrix multiplication problem in semirings. We assume that initially, matrices $A$ and $B$ are distributed evenly among all the processors. The structure of the distribution does not matter because we can reach any desired even distribution in one round (even if we start with an uneven distribution). At the end of the algorithm the elements of the output matrix $C$ should also be evenly distributed. We analyse the problem in the following three settings.
  1. 1.
    When $A$ and $B$ are square matrices of size $n\times n$. We call this $(n,n,n)$ case.
  2. 2.
    $A$ and $B$ are rectangular matrices. We divide it into the following two cases
    1. (a)
      $(n,d,n):A$ and $B$ are of size $n\times d$ and $d\times n$ respectively
    2. (b)
      $(d,n,d):A$ and $B$ are of size $d\times n$ and $n\times d$ respectively
  3. 3.
    $A$ and $B$ are $d$-sparse $n\times n$ matrices, i.e. each row and column of matrices $A$ and $B$ contains at most $d$ non-zero elements. We call it the $d$-sparse case.
In this paper, we also assume that $P.M\leq O(N)$ i.e. the total memory of the system is just tight enough to hold the input and output matrices.
1.3 Our Results
We prove that $(n,d,n)$ case can be solved in $O(n^{\frac{\alpha}{2}})$ rounds. Our algorithm that gives this bound is very straightforward. However, what is interesting is that we prove that this simple algorithm is the best we can hope for, i.e., we prove the matching lower bound. We prove that it will take $\Omega(n^{\frac{\alpha}{2}})$ rounds to solve the problem. For $(n,d,n)$ case, we prove that the problem can be solved in $\Theta(\frac{d}{\sqrt{n}})$ rounds with $n$ processors each with $O(n)$ memory. Similarly, for the $(d,n,d)$ case, we prove that it will take $\Theta(\sqrt{d}+\log_{d}n)$ round when number of processors are $n$, and $\Theta(\frac{d}{\sqrt{n}})$ rounds when number of processors are $d$. For the $d$-sparse case, when the number of processors is $n$, we improve the trivial bound of $O(d)$ to $O(d^{0.9})$. In this case, we could not prove a lower bound, though we notice that some instances of this type can be solved in $O(\sqrt{d})$ rounds (when all the elements are concentrated along the diagonal in the input matrices).
1.3.1 High-level Idea of Our Technique
For square and rectangular cases, we observed that in order to reduce inter-processor communication, the best case is when all processors try to compute a square submatrix of $C$. Thus, we divide the matrices $A$ and $B$ into smaller square matrices of equal size depending on the local memory of the processor, and assign each submatrix to one of the processors. At the end, each processor needs to produce the corresponding submatrix of $C$. The interesting part lies in the lower bound. Since we are working with semirings, we know that we need to compute a total $n^{3}$ products (in case of $(n,n,n)$, and $n^{2}d$ and $d^{2}n$ products in case of $(n,d,n)$ and $(d,n,d)$ respectively) of type $a_{ik}b_{kj}$, where $a_{ik}$ and $b_{kj}$ are the elements of matrices $A$ and $B$ respectively. We analyse the maximum number of products (we call it a term in this paper) a processor can compute in each round. Since we are working with semirings, in order to compute an element of matrix $C$, a processor needs to have information about all the terms required to compute that element explicitly. We show that in $(n,n,n)$ case, because of the memory restriction of $O(n^{2-\alpha})$, a processor cannot compute more than $O(n^{\frac{3}{2}\alpha})$ terms in one round. Since there are $n^{\alpha}$ processors, it will take $\Omega(n^{\frac{\alpha}{2}})$ rounds. A similar argument works for rectangular cases.
For the $d$-sparse case, we extend the idea of [4]. In their paper, they show that $d$-sparse matrix multiplication is equivalent to processing triangles in a $d$-degree graph. In other words, processing triangles is equivalent to computing terms. They show that given any $d$-degree graph, say $T$, we can divide the graph into two parts such that one part, say $T^{\prime}$, is clustered - it contains layers of subgraphs of the original graph such that in each layer we have disjoint clusters of triangles which are small and dense, and the other part, say $T^{\prime}$, contains a low number of triangles. Now, in order to process all the triangles, we can first process all the triangles in the clustered part layer-by-layer using square matrix multiplication, then compute all the triangles $T^{\prime}$ using brute force. This is a graph-theoretical result independent of any model. In this paper, we refined their result (we will describe it in a later section) and then use our square matrix multiplication as a black box to solve the clustered part.
2 Square Matrix Multiplication
In this section, we will prove that if there are $n^{\alpha}$ processors each with $O(n^{2-\alpha})$ memory, then $(n,n,n)$ case can be solved in $O(n^{\alpha/2})$ rounds in the MPC model. This required memory is tight, as the size of input and output is $O(n^{2})$. For the sake of simplicity, we first give the analysis of the case when $\alpha=1$. Then we prove the result in full generality.
Theorem 2.1.
Matrix multiplication in MPC model with $n$ processors, each with $O(n)$ memory, can be performed in $O(\sqrt{n})$ rounds for semirings.
Proof 2.2.
We shall prove this by proposing a simple algorithm which achieves $O(\sqrt{n})$ bound. Divide each of the matrices $A$, $B$ and $C$, into $n$ sub-matrices of size $\sqrt{n}\times\sqrt{n}$. Let $M^{ij}$ denote the sub-matrix which consists of elements of $M$ at rows $(i-1)\sqrt{n}+1,(i-1)\sqrt{n}+2,\ldots,i\sqrt{n}$ and columns $(j-1)\sqrt{n}+1,(j-1)\sqrt{n}+2,\ldots,j\sqrt{n}$ (where $M\in\{A,B,C\})$ Let us index the $n$ processors as $p_{11},p_{12},\ldots,p_{\sqrt{n}\sqrt{n}}$. Assign the submatrices $A^{ij}$, $B^{ij}$ and $C^{ij}$ to processor $p_{ij}$. Initialize $C^{ij}$ as a null matrix at $p_{ij},\forall\ i,\ j$. The algorithm works as follows.
for $x\leftarrow 1$ to $\sqrt{n}$ do
       send all elements of $A^{ix}$ and $B^{xj}$ to $p^{ij}$ $\forall\ i,\ j\leq\sqrt{n}$
$p_{ij}$ locally computes $C^{ij}$ = $C^{ij}$ + $A^{ix}*B^{xj}$ $\forall\ i,\ j$ and * denotes matrix multiplication
(Note that this step can be done without exceeding $O(n)$ memory)
Algorithm 1 Square Matrix Multiplication in $O(\sqrt{n})$ rounds
The above algorithm follows all the constraints of the MPC model. It is easy to see that during the course of the algorithm, each node sends at most $n\sqrt{n}$ elements to other nodes and receives at most $n\sqrt{n}$ elements from other nodes. Since each element can send and receive $O(n)$ elements in each round, the above algorithm will terminate in $O(\sqrt{n})$ rounds. At the end, processor $p_{ij}$ will contain the submatrix $C^{ij}$ of the final output matrix $C$. This finishes the proof of Theorem 2.1.
Now we will proceed towards proving the lower bound and prove that the simple algorithm presented above is actually tight and cannot be improved further. Before that, we will prove the following lemma, which we will use to prove all the lower bounds that we present in this paper.
Lemma 2.3.
A processor with $r$ memory can calculate at most $O(r\sqrt{r})$ terms in one round in the MPC model for any matrix multiplication instance.
Proof 2.4.
We prove the lower bound by finding the limit on computational capacity per processor. We know that $c_{ij}$ = $\sum_{x=1}^{n}a_{ix}*b_{xj},$ $\forall\ i,\ j$. As mentioned earlier, we refer to $a_{ix}*b_{xj}$ as one term of matrix $C$ (since it is used to compute an element of $C$). We try to find the maximum number of terms that a processor can calculate in one round. Let us assume that $t$ is the maximum number of terms of $C$ that a processor can calculate in one round. Each of these terms is of the form $a_{ix}*b_{xj}$. We segregate the terms into groups on the basis of the values of $x$, so that the terms with the same value of $x$ are in the same group. Let $H_{1},H_{2},\ldots$ be different groups containing $t_{1},t_{2},t_{3},...$ terms respectively (see Figure 1). We try to find the minimum memory required by a processor to calculate these terms. We know that in the case of semirings, in order to compute a term, say $a_{ix}*b_{xj}$, a processor needs to receive both elements $a_{ix}$ and $b_{xj}$ distinctly. Notice that no elements of $A$ and $B$ are common between the terms present in any two different groups. So we try to find the minimum memory required to store the elements of $A$ and $B$ involved in the terms of group $H_{i}$, and sum it for all values of $i$. This sum should not exceed the $O(n)$ memory available with a processor. Let $p$ and $q$ be the number of elements of $A$ and $B$ required to calculate the terms of $H_{i}$. We can say that, $p*q\geq t_{i}$ (since each term is unique, $t_{i}$ combinations should exist). This implies that $p+q\geq 2\sqrt{t_{i}}$. Thus, at least $2\sqrt{t_{i}}$ memory is required per group, leading to the total memory requirement of $2\sum\sqrt{t_{i}}\leq O(r)$. Notice that the processor computing terms of $H_{k}$ needs to keep the value of all the terms (i.e the result of the product $a_{ik}b_{kj})$ present in $H_{i}$ in different memory locations. Because any two terms, say $a_{i_{1}k}*b_{kj_{1}}$ and $a_{i_{2}k}*b_{kj_{2}}$ present in $H_{i}$, will appear in the computation of $c_{i_{1}j_{1}}$ and $c_{i_{2}j_{2}}$. Thus, the processor cannot mix them in any way (since we are working with semirings). Therefore, the number of terms in a group should not exceed the memory of the processor. From this, we can conclude that
$t_{i}\leq O(r),\ \forall\ i$
and from above, we know that
$\sum\sqrt{t_{i}}\leq O(r)$
Thus, $\sum t_{i}=t$ is maximized when $t_{i}=\Theta(r)$ and the number of groups = $\Theta(\sqrt{r})$. Therefore, $t=O(n\sqrt{n})$ is the maximum number of terms that a processor can calculate in one round.
Figure 1: $n^{3}$ terms of matrix $C$ represented in the form of $n$ layers (each layer of distinct colour). One layer is obtained by multiplying one column of $A$ with one row of $B$. Each $H_{i}$ consists of terms only from one of the layers (here, three red filled cubes from Layer 1).
Theorem 2.5.
Matrix Multiplication in the MPC model with $n$ processors, each with $O(n)$ memory, requires $\Omega(\sqrt{n})$ rounds in semirings.
Proof 2.6.
As there are $n$ processors with $O(n)$ memory, a maximum of $\Theta(n^{2}\sqrt{n})$ terms can be calculated in one round. There are a total of $n^{3}$ terms in $C$. Hence, we need at least $\Omega(\sqrt{n})$ rounds to calculate the product matrix $C$.
From Theorem 2.1 and Theorem 2.5, we can conclude the following.
Theorem 2.7.
Square matrix multiplication in the MPC model with $n$ processors, each with $O(n)$ memory, requires $\Theta(\sqrt{n})$ rounds in semirings.
Let’s analyse the generalised case now. We have $n^{\alpha}$ processors where $0\leq\alpha\leq 2$. To keep the MPC model tight, each processor has $O(n^{2-\alpha})$ memory, so that the total memory is $O(n^{2})$.
Theorem 2.8.
Matrix multiplication in MPC model with $n^{\alpha}$ processors each with $O(n^{2-\alpha})$ memory, takes $\Theta(n^{\alpha/2})$ rounds for semirings.
Proof 2.9.
Similar to Theorem 2.1, we divide the matrices $A$, $B$, and $C$ into $n^{\alpha}$ sub-matrices of size $n^{1-\frac{\alpha}{2}}\times n^{1-\frac{\alpha}{2}}$. Let $M^{ij}$ and $p_{ij}$ be defined and distributed in the same way, except $i,j\in[n^{\frac{\alpha}{2}}]$ (instead of $\sqrt{n}$). We run the following algorithm in order to compute the matrix $C$.
for each $x$ from $1$ to $n^{\alpha/2}$ do
       Send elements of $A^{ix}$ and $B^{xj}$ to $p_{ij}$ $\forall\ i,\ j$
$p_{ij}$ locally computes $C^{ij}$ = $C^{ij}$ + $A^{ix}*B^{xj}$ $\forall\ i,\ j$
(* denotes Matrix Multiplication)
(Note that this step can be done without exceeding $O(n^{2-\alpha})$ memory)
Algorithm 2 Square Matrix Multiplication in $O(n^{\alpha/2})$ rounds
The above algorithm follows all the constraints of the MPC model. At the end of the algorithm, $p_{ij}$ contains $C^{ij}$ for all $1\leq i,j\leq n^{\alpha/2}$, meaning the entire product matrix $C$ has been calculated. Observing the argument presented in the proof of Theorem 2.5, in this setting, we can say that a processor can calculate at most $O(n^{3-3\alpha/2})$ terms in one round. Thus, a maximum of $O(n^{3-\alpha/2})$ terms can be calculated in total in one round. Hence, we need at least $\Omega(n^{\alpha/2})$ rounds to calculate the product matrix $C$. So, the given algorithm is the best possible algorithm for square matrix multiplication of general semirings in the MPC model.
3 Rectangular Matrix Multiplication
In this section, we consider the case when matrices $A$ and $B$ are rectangular matrices. There can be many types of instances in rectangular matrix multiplication. In this paper, we focus on two primary cases of type $(n,d,n)$ and $(d,n,d)$ and analyse them separately.
3.1 (n,d,n) Matrix Multiplicaton
In this case, we assume that we have $n$ processors with $O(n)$ memory each. Notice that in this case the matrices $A$ and $B$ can be stored in $O(nd)$ memory; however, the resulting $C$ matrix is a square matrix of size $n\times n$. Thus, we limit the total memory of the system to $O(n^{2})$. We divide the matrix $C$ into $n$ sub-matrices of size $\sqrt{n}\times\sqrt{n}$ each, denoted by $C^{ij}$ (similar to Section 2) where $i,j\in[\sqrt{n}]$. Assign the computation of $C^{ij}$ to processor $p^{ij}$. Without loss of generality, assume that we scale the value of $d$ such that $d:=(\lceil\frac{d}{\sqrt{n}}\rceil\sqrt{n}).$ Divide the matrices $A$ and $B$ into sub-matrices of size $\sqrt{n}\times\sqrt{n}$, each being denoted as $A^{iq}$ and $B^{qj}$ respectively, where $i,\ j\in[\sqrt{n}]$ and $q\in[\frac{d}{\sqrt{n}}]$. Observe that
$C^{ij}=\sum_{q=1}^{\frac{d}{\sqrt{n}}}A^{iq}*B^{qj},\text{where $*$ denotes matrix multiplication}$
We can assume that each processor $p_{ij}$ contains one row and one column of $A$ and $B$ respectively . Note that an element of $A$ needs to be sent to at most $\sqrt{n}$ processors. Thus, $p_{ij}$ needs to send $d\sqrt{n}$ elements. In one round, it can send $n$ elements and thus, in $\frac{d}{\sqrt{n}}$ rounds, it can send all the elements to the destination processors. Now let us analyse how many elements of $A$ and $B$ are received by each processor. $A^{iq}$ and $B^{qj}$ are $\sqrt{n}\times\sqrt{n}$ matrices containing a total of $n$ elements. Therefore, $p_{ij}$ needs a total of $(\frac{d}{\sqrt{n}}\cdot n)=d\sqrt{n}$ elements of $A$ and $B$ in order to compute $C^{ij}$. In one round, $p_{ij}$ can receive at most $O(n)$ elements and thus, in $\frac{d\sqrt{n}}{n}=\frac{d}{\sqrt{n}}$ rounds, it can receive all the elements that it requires for the computation of $C^{ij}$. Hence, we get the following simple algorithm.
for each $q$ from $1$ to $\frac{d}{\sqrt{n}}$ do
       receive elements of $A^{iq}$ and $B^{qj}$ from processors containing them
$p_{ij}$ locally computes $C^{ij}$ = $A^{iq}*B^{qj}$
Algorithm 3 $(n,d,n)$ Matrix Multiplication
The lower bound follows from Lemma 2.3. There are $n^{2}d$ terms that need to be calculated. In one round, a total of $n^{2}\sqrt{n}$ terms can be calculated. Therefore, $\Omega(\frac{d}{\sqrt{n}})$ rounds will be required. We proved the following theorem.
Theorem 3.1.
Multiplication of two rectangular matrices of size $n\times d$ and $d\times n$ respectively, with $n$ processors of $O(n)$ memory requires $\Theta(\frac{d}{\sqrt{n}})$ rounds in semirings, where $d\leq n$.
3.2 (d,n,d) Matrix Multiplication
In the $(d,n,d)$ case, the size of all three matrices is bounded by $nd$; thus, it is sufficient to have a total memory of size $O(nd)$. We further divide this case into two subcases: (i) when there are $n$ processors each with $O(d)$ memory, and (ii) when there are $d$ processors each with $O(n)$. First, we prove the following two lemmas that will help us to analyse both these cases.
Figure 2: Upper layer represents the initial distribution of elements, and lower layer represents the distribution after the first round
Lemma 3.2.
If there are $t$ processors each containing matrices of size $\sqrt{k}\times\sqrt{k}$, and the memory of each processor is $O(k)$, then the sum of these matrices can be computed in $O(\log_{k}t)$ rounds in the MPC model.
Proof 3.3.
Suppose $M_{1},M_{2},\ldots,M_{t}$ are the $t$ matrices and $X=\sum M_{l}$. Let $m_{ij}^{l}$ denote the element at $i^{th}$ row and $j^{th}$ column of $M_{l}$. Therefore $x_{ij}=\sum m_{ij}^{l}$. We divide the $t$ processors into $k$ groups of $(\frac{t}{k})$ processors such that each group is responsible for computing one $x_{ij}$. Let’s just focus on the first group, the same algorithm will be followed in the other groups. Let $p_{1},p_{2},\ldots,p_{t/k}$ be the processors in the first group. In the first round, all $t$ processor exchange messages such that $p_{1}$ receives elements $m_{11}^{1},m_{11}^{2},\ldots m_{11}^{k}$; $p_{2}$ receives elements $m_{11}^{k+1},m_{11}^{k+2},\ldots m_{11}^{2k}$ and so on (see Figure 2 ). Now, each $p_{i}$ computes the sum of all the elements it receives. Notice that all the values that are required to compute $x_{11}$ are present in $t/k$ processors, such that each processor contains only one value (the sum it has calculated). Now, in the next round, processor $p_{2},p_{3},\ldots p_{k}$ will send all the intermediate values they have calculated in the previous round to $p_{1}$. Similarly, $p_{k+2},p_{k+3},\ldots p_{2k}$ will send all the intermediate values they have calculated in the previous round to $p_{k+1}$ and so on (see Figure 3for reference). Now, notice that all the values that are required to compute $x_{11}$ are now present in the $(\frac{n}{k^{2}})$ processors, and each contains only one value. Repeat the previous step for $O(\log_{k}t)$ rounds. After $O(\log_{k}t)$, the final value of $x_{11}$ will be present in one processor. The same thing will happen in the other groups. Therefore, after $O(\log_{k}t)$ rounds processors will calculate $X$. This finishes the proof of the lemma. Notice that in the case where $t\leq k$, the algorithm will terminate in merely a single round. Therefore, in the above proof, we can assume $t>k$.
Figure 3: Processors in the first group computing the value of $x_{11}$
Lemma 3.4.
If there are $t$ processors each with $O(k)$ memory such that each processor contains one element, then $\Omega(\log_{k}t)$ rounds are required to compute the sum of these elements in the MPC model, for semirings.
Proof 3.5.
To compute the sum, the only thing we can do is add the elements locally during a round of internal computation, and then send or receive this local sum(s) in some fashion during a communication round. We try to follow $x$–the number of processors that hold elements yet to be added together to obtain the final sum. Initially, $x$ is equal to $t$, and we need to reduce it to one in order to add all the elements. In one communication round, this $x$ can never decrease by more than a factor of $O(k)$, as one processor can only receive information from $O(k)$ processors during a communication round, due to its memory constraint. Thus, to reduce $x$ to one (which is the same as adding all elements), we will require $\Omega(\log_{k}t)$ communication rounds.
Now we are all set to analyse the (d,n,d) case in the aforementioned settings. First, we discuss the case with $n$ processors with $O(d)$ memory.
3.2.1 n Processors with O(d) Memory
Let us divide the matrix $C$ into $d$ sub-matrices of size $\sqrt{d}\times\sqrt{d}$ each, denoted by $C^{ij}$, where $i,j\in[\sqrt{d}]$. We assign $(\frac{n}{d})$ processors to evaluate each sub-matrix. Also divide the matrices $A$ and $B$ into sub-matrices of size $\sqrt{d}\times\sqrt{d}$, each being denoted as $A^{iq}$ and $B^{qj}$ respectively, where $i,j\in[\sqrt{d}]$ and $q\in[\frac{n}{\sqrt{d}}]$. We know that,
$C^{ij}=\sum_{q}A^{iq}*B^{qj}$
Observe that there would be a total of $n$ submatrices of the forms $A^{iq}$ and $B^{qj}$. In one round of communication, let’s distribute the input among the $n$ processors such that each processor contains exactly two sub-matrices, namely, $A^{iq}$ and $B^{qj}$ for some values of $i,j,q$ such that the sub-matrices stored by any two processors are distinct. This is possible since storing $A^{iq}$ and $B^{qj}$ requires only $O(d)$ memory. Let us focus on computation on $C^{11}$, other submatrices are computed in the same way. We have assigned $\frac{n}{d}$ processors to compute it. Let $P^{ij}=\{p_{1},p_{2},\ldots p_{\frac{n}{d}}\}$ be the set of processor assigned to $C^{ij}$. In round $k$, processor $p_{i}$ receives submatrices corresponding to $q=(i-1)\sqrt{d}+k$ i.e. in round 1, it receives $A^{1((i-1)\sqrt{d}+1)}$ and $B^{((i-1)\sqrt{d}+1)1}$, in round 2, it receives $A^{1((i-1)\sqrt{d}+2)}$ and $B^{((i-1)\sqrt{d}+2)1}$ and so on. Processor $p_{i}$ then computes the product of these submatrices locally after each round, adding the newly computed product to the existing product that it computed in the previous round using simple matrix addition. It’s easy to observe that after $\sqrt{d}$ rounds, processor $p_{i}$ will contain $\sum_{k=1}^{\sqrt{d}}A^{1((i-1)\sqrt{d}+k)}*B^{((i-1)\sqrt{d}+k)1}$. Now we have $(\frac{n}{d})$ submatrices distributed over the processors in $P^{ij}$, each containing a matrix of size $\sqrt{d}\times\sqrt{d}$ such that the sum of those matrices will produce $C^{11}$. According to Lemma 3.2, $C^{11}$ can be computed in $O(\log_{d}n)$ rounds. Similarly, the other $C^{ij}$ are also computed in parallel. Thus, matrix $C$ can be computed in $O(\sqrt{d}+\log_{d}n)$ rounds. The only thing missing in the proof is to show that each processor is sending messages of $O(d)$ size per round. Notice that each element of matrices $A$ and $B$ needs to be sent to at most $\sqrt{d}$ processors. Each processor initially holds $d$ elements of both $A$ and $B$. Thus, it needs to send $d\sqrt{d}$ elements to another processor, which can easily be done in $O(\sqrt{d})$ rounds, such that each processor sends $O(d)$ elements in each round. This finishes the proof for the upper bound.
Lower Bound:
The lower bound follows from Lemma 2.3 and Lemma 3.4. There are $d^{2}n$ terms that need to be calculated. In one round, a total of $O(nd\sqrt{d})$ terms can be calculated. Therefore, $\Omega(\sqrt{d})$ rounds will be required. We know that $c_{ij}=\Sigma_{k=1}^{n}a_{ik}b_{kj}$. Without loss of generality assume that all $b_{kj}=1$. Therefore, the processor computing $c_{ij}$ needs to compute the sum of $n$ elements that are initially present in at least $\frac{n}{d}$ processors because of $O(d)$ memory restriction. Hence, it will require $\Omega(\log_{d}n)$ rounds. Therefore, the final lower bound is $\Omega(\sqrt{d}+\log_{d}n)$
Theorem 3.6.
Multiplication of two rectangular matrices of size $d\times n$ and $n\times d$ respectively, with $n$ processors of $O(d)$ memory requires $\Theta(\sqrt{d}+\log_{d}n)$ rounds in semirings, where $d\leq n$.
3.2.2 d processors with O(n) memory
We handle this case also in a similar way to the previous one, except the division of processors and matrices is slightly different. Without loss of generality, assume that we scale the value of $d$ such that $d:=(\lceil\frac{d}{\sqrt{n}}\rceil\sqrt{n})$. Let us divide the matrix $C$ into $(\frac{d^{2}}{n})$ sub-matrices of size $\sqrt{n}\times\sqrt{n}$ each, denoted by $C^{ij}$, where $i,j\in[\frac{d}{\sqrt{n}}]$. We assign $(\frac{n}{d})$ processors to evaluate each sub-matrix. Also divide the matrices $A$ and $B$ into sub-matrices of size $\sqrt{n}\times\sqrt{n}$, each being denoted as $A^{iq}$ and $B^{qj}$ respectively, where $i,j\in[\frac{d}{\sqrt{n}}]$ and $q\in[\sqrt{n}]$. We know that,
$C^{ij}=\sum_{q}A^{iq}*B^{qj}$
Observe that there would be a total of $d$ submatrices of the form $A^{iq}$ and $B^{qj}$. In one round of communication, let’s distribute the input among the $d$ processors such that each processor contains exactly two sub-matrices, namely, $A^{iq}$ and $B^{qj}$ for some values of $i,j,q$ such that the sub-matrices stored by any two processors are distinct. This is possible since storing $A^{iq}$ and $B^{qj}$ requires only $O(n)$ memory. Let us focus on computation on $C^{11}$, other submatrices are computed in the same way. We have assigned $\frac{n}{d}$ processors to compute it. Let $P^{ij}=\{p_{1},p_{2},\ldots p_{\frac{n}{d}}\}$. In round $k$, processor $p_{i}$ receives submatrices corresponding to $q=(i-1)\frac{d}{\sqrt{n}}+k$ i.e. in round 1, it receives $A^{1((i-1)\frac{d}{\sqrt{n}}+1)}$ and $B^{((i-1)\frac{d}{\sqrt{n}}+1)1}$, in round 2, it receives $A^{1((i-1)\frac{d}{\sqrt{n}}+2)}$ and $B^{((i-1)\frac{d}{\sqrt{n}}+2)1}$ and so on. It is easy to observe that $p_{i}$ would receive these sub-matrices from a single other processor because of the way in which the initial distribution of $A^{iq}$ and $B^{qj}$ was done among the processors. Thus, the amount of information that a processor ends up sending in any communication round would be $O(n)$. Processor $p_{i}$ then computes the product of these submatrices locally after each round, adding the newly computed product to the existing product that it computed in the previous round using simple matrix addition. It’s easy to observe that after $\frac{d}{\sqrt{n}}$ rounds, processor $p_{i}$ will contain $\sum_{k=1}^{\frac{d}{\sqrt{n}}}A^{1((i-1)\frac{d}{\sqrt{n}}+k)}*B^{((i-1)\frac{d}{\sqrt{n}}+k)1}$. Now we have $(\frac{n}{d})$ submatrices distributed over the processors in $P^{ij}$, each containing a matrix of size $\sqrt{n}\times\sqrt{n}$ such that the sum of those matrices will produce $C^{11}$. According to Lemma 3.2, $C^{11}$ can be computed in $O(\log_{n}\frac{n}{d})$ rounds. But since $\frac{n}{d}\leq n$, it can be computed in a constant number of rounds. Similarly, the other $C^{ij}$ are also computed in parallel. Thus, matrix $C$ can be computed in $O(\frac{d}{\sqrt{n}})$ rounds. The lower bound follows from Lemma 2.3. There are $d^{2}n$ terms that need to be calculated. In one round, a total of $O(dn\sqrt{n})$ terms can be calculated. Therefore, $\Omega(\frac{d}{\sqrt{n}})$ rounds will be required. We proved the following theorem.
Theorem 3.7.
Multiplication of two rectangular matrices of size $d\times n$ and $n\times d$ respectively, with $d$ processors of $O(n)$ memory requires $\Theta(\frac{d}{\sqrt{n}})$ rounds in semirings, where $d\leq n$.
4 Sparse Matrix Multiplication
In this section, we will discuss the case when $A$ and $B$ are $d$-sparse. Notice that in this case, the resulting matrix $C=A*B$ can be $d^{2}$-sparse. But in the same spirit as discussed in Gupta et al. [4], in order to keep the model more meaningful, we assume that we are interested in only at most $d$ elements of each row and column of $C$. Therefore, both the input and output can be stored on $O(nd)$ memory. We analyse the case when we have $O(n)$ processors, each with $O(d)$ memory. We assume that initially each processor holds one row and one column (i.e. $O(d)$ elements in total) of $A$ and $B$, and at the end each processor outputs $d$ elements of one of the rows of $C$. We assume that the position of these $d$ elements of $C$ that a processor needs to output is already known to the processor.
A trivial algorithm:
Notice each element of $C$ is the sum of $d$ terms. Thus, there are $nd^{2}$ terms that need to be calculated. In one round, a processor can compute any $d$ terms of $C$. It can do this by obtaining the $d$ elements of matrices $A$ and $B$ which are present in the terms, and then performing $d$ multiplications to compute the $d$ terms. Thus, in one round of local computation, the $n$ processors can compute any $nd$ terms of $C$. Hence, to compute the $nd^{2}$ of $C$, we only require $O(d)$ communication rounds. Now, we give a better algorithm than this as follows.
We try to find a more efficient method to compute the terms of $C$. For that, we draw inspiration for Gupta et al. [4]. In particular, we use Lemma 4.2 of their paper. As we discussed earlier, in their paper, they translated the problem of matrix multiplication to triangle processing; here we try to restate their results according to the terminology we used in this paper as follows.
Lemma 4.1 ([4]).
Let $0\leq\varepsilon_{1}<\varepsilon_{2}$, and assume $d$ is sufficiently large. Let $T$ denote the remaining number of terms of $C$ that need to be computed such that
$|T|=O(d^{2-\epsilon_{1}}n).$
Then, we can compute these terms in $l$ sequential iterations where
$l=O(d^{5\epsilon_{2}-\epsilon_{1}})$
and after these iterations only $T^{\prime}$ terms remain to be computed, such that
$|T^{\prime}|=O(d^{2-\epsilon_{2}}n)$
Let us try to explain the lemma in simple words. The lemma basically states that the multiplication of two $n\times n$, $d$-sparse matrices ($C=A*B$) that contain $O(d^{2-\epsilon_{1}}n)$ terms to be computed can be divided into two parts $C^{\prime}$ and $C^{\prime\prime}$ such that
  • $C^{\prime}=A_{1}\times B_{1}+A_{2}\times B_{2}+\ldots+A_{l}\times B_{l}$ ($l$ is same as mentioned in the lemma) where $C_{i}=A_{i}\times B_{i}$ such that both $A_{i}$ and $B_{i}$ consist of $(\frac{n}{d})$ submatrices $A_{i}^{j}$ and $B_{i}^{j}$, for $j\in[\frac{n}{d}]$, of size $d\times d$ such that $A_{i}\times B_{i}=\sum A_{i}^{j}*B_{i}^{j}$
  • $C^{\prime\prime}=A^{\prime}\times B^{\prime}$ such that, it requires at most $O(d^{2-\epsilon_{2}}n)$ terms to compute $C^{\prime\prime}$
Thus, the total time required will be the time to compute $C^{\prime}$ and $C^{\prime\prime}$. The idea is to use square matrix multiplication as a black box algorithm to compute $C^{\prime}$ and use a brute force algorithm to compute $C^{\prime\prime}$. To compute $C^{\prime}$, we can run $l$ sequential iterations, each for one $A_{i}\times B_{i}$. In one iteration, we can assign $d$ processors for the computation of $A_{i}^{j}*B_{i}^{j}$ for each $j$. These $(\frac{n}{d})$ computations will happen in parallel in one iteration. In this paper, we improve the bound on the number of iterations ($l$) required to go from $T$ to $T^{\prime}$. Notice that, in the MPC model, the time required for one iteration is $O(\sqrt{d})$.
4.1 Tight Bound for number of Iterations
We prove a tight upper bound on the number of iterations required to reach from $T$ to $T^{\prime}$(mentioned in Lemma 4.1).
Lemma 4.2.
Let $0\leq\varepsilon_{1}<\varepsilon_{2}$. Suppose $|T|\;=\;O(d^{2-\varepsilon_{1}}n)\quad\text{and}\quad|T^{\prime}|\;=\;O(d^{2-\varepsilon_{2}}n)$. Then the total number of iterations needed to reduce the number of uncomputed terms from $T$ to $T^{\prime}$ is $l\;=\;O\!\bigl{(}d^{4\,\varepsilon_{2}}\bigr{)}$.
Proof 4.3.
Let’s observe what happens when we introduce an intermediate step in the reduction of terms from $T$ to $T^{\prime}$. Let $0\leq\epsilon_{1}<\epsilon_{3}<\epsilon_{2}$ such that $\varepsilon_{3}\;=\;\frac{\varepsilon_{1}+\varepsilon_{2}}{2}$. Suppose we have
$|T|\;=\;O(d^{2-\varepsilon_{1}}n)\quad\text{,}\quad|T^{\prime\prime}|\;=\;O(d^{2-\varepsilon_{3}}n)\quad\text{, and }\quad|T^{\prime}|\;=\;O(d^{2-\varepsilon_{2}}n).$
Let $l_{1}$ denote the number of iterations required to reduce the number of uncomputed terms from $T$ to $T^{\prime\prime}$, and similarly, $l_{2}$ denote the number of iterations required to reduce the number of uncomputed terms from $T^{\prime\prime}$ to $T^{\prime}$. By Lemma 4.1, we have $l_{1}=O(d^{5\epsilon_{3}-\epsilon_{1}})$ and $l_{2}=O(d^{5\epsilon_{2}-\epsilon_{3}})$. Thus, the total number of iterations required to reduce the number of uncomputed terms from $T$ to $T^{\prime}$ is $l=l_{1}+l_{2}=O(d^{5\epsilon_{3}-\epsilon_{1}})+O(d^{5\epsilon_{2}-\epsilon_{3}})=O(d^{5\epsilon_{2}-\epsilon_{3}})$, (since $5\epsilon_{2}-\epsilon_{3}>{5\epsilon_{3}-\epsilon_{1}}$).
Now, if we recursively insert $\epsilon_{4}$ between $\epsilon_{3}$ and $\epsilon_{2}$ (such that $\epsilon_{4}$ is the arithmetic mean of $\epsilon_{3}$ and $\epsilon_{2}$) and similarly $\epsilon_{5}$ between $\epsilon_{4}$ and $\epsilon_{2}$ and do this $x$ number of times ($x$ is finite), then, we can see that number of iterations required to reduce the number of terms step by step would be $l=l_{1}+l_{2}+..+l_{x+1}$, where $l_{1}$ is the number of iterations to go from $\epsilon_{1}$ to $\epsilon_{3}$, $l_{2}$ from $\epsilon_{3}$ to $\epsilon_{4}$, .., $l_{x+1}$ from $\epsilon_{x+2}$ to $\epsilon_{2}$. Observe that the order of $l$ is dominated by the term $l_{x+1}$. Thus, $l=O(d^{5\epsilon_{2}-\epsilon_{x+2}})$. Hence, as $x$ becomes very large, $l$ converges to
$O(d^{5\epsilon_{2}-\epsilon_{x+2}})\approx O(d^{4\epsilon_{2}})$
Final Computation of terms:
As we mentioned, the multiplication of two $d$-sparse matrices can be done in two phases. In one phase, we iteratively use square matrix multiplication, and in the other phase, we use the trivial algorithm. Our task is to compute $O(d^{2}n)$ terms. We will first reduce the number of uncomputed terms from $O(nd^{2})$ to $O(nd^{2-\epsilon})$ ($\epsilon\geq 0$) by iteratively using square matrix multiplication, and then compute the rest of the terms using trivial computation. By Lemma 4.2, the number of iterations to reduce the number of uncomputed terms from $O(nd^{2})$ to $O(nd^{2-\epsilon})$ is $O(d^{4\epsilon})$ where each iteration takes $O(\sqrt{d})$ communication rounds (complexity of square matrix multiplication). And computing $O(nd^{2-\epsilon})$ terms using trivial computation takes $O(\frac{nd^{2-\epsilon}}{nd})=O(d^{1-\epsilon})$communication rounds.Thus, the total number of communication rounds needed is $O(d^{4\epsilon+1/2})+O(d^{1-\epsilon})$, which is minimized at $\epsilon=0.1$ taking $O(d^{0.9})$ communication rounds. Thus, we prove the following theorem
Theorem 4.4.
Multiplication of two $d$-sparse matrices in the MPC model with $n$ processors, each with $O(d)$ memory, can be performed in $O(d^{0.9})$ rounds in semirings.